Recent Log Entries

CREW CHIEF SIGN IN

Log Number: 3819129Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 08-01-20
Time : 23:40:09
Day : Saturday
Shift : Owl

Crew Chief : Merz
Program Deputy: Benesch
Team : D
Arm : McCaughan
Designated SSO: Anthony
LERF Operator:
Operators : Anthony, McCaughan

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

Follow-up Re: building temps

Log Number: 3819112Logbook(s): Body: 

Every bit of that 88 degrees at zone 1L14 at noon to 1500. It felt like 90 percent humidity and causing electrical havoc on the contactors and electrical connections. Sweat was running on my glasses inside and fogging outside !

Follow-up Re: 0L03-4 Tuner Issues

Log Number: 3819110Logbook(s): Body: 

If you see GASK(Mod Drive) railing at 9.5 v or back and forth and the TDETA all over, that could be several possibilities. A loss of RF drive from the RFCM, the loss or decline of the 2.5w amp, or the cavity can't find the tune frequency due to vacuum or cryo instability. We RF experts have to examine the expert pages and graph all the other possible causes outside of hardware so we need to be called. I missed the call from S. Holben due to a loss of power due to the storm and my internet outage and apologize ! These storms are messing with a lot more than the C100 stability.

0L03-4 TDETA at Time 1548

Log Number: 3819105Logbook(s): Entry Makers: kernsBody: 

Cavity 3/4 failed at 2 and 3 minutes after the vacuum spike at 1548. TDETA will go bonkers when then cavity is trying to tune with bad vacuum and especially when the gradient is set above what the cavity was previously running. Cavity 4 was brought back at 8.979 Mv/m instead of 6 Mv/m which took both cavity out again on CWVF. The amp and RFCM looked ok before the vacuum fault and any time the cavities TDETA are jumping around is when vacuum is upset or gradients are higher than what they need to prevent CWVF faults.

1L19-1/2 Vacuum Problems

Log Number: 3819102Logbook(s): Entry Makers: kernsBody: 

Cavity 1 RFCM was replaced yesterday ~noon after TDETA jumping. Cavity 2 was replaced the day before for the same thing. Vacuum spiked yesterday and took cavity 1 and 2 out but was blamed on the RFCM and not the bad vacuum when cavity 1 was put at 13Mv/m vice the ODVH=6.7 Mv/m. and cavity 2 at ODVH=8.5 Mv/m. Cavity 2 has since been lowered to 5.886 Mv/m presumably to reduce trips do to vacuum as indicated on the plot. Cavity 1 and 2 was still being affected by CWVF at approx.

Autolog save

Log Number: 3819098Logbook(s): Tags: Body: saver ran from opsl632

Saved: eforman 08-01-20 21:34:32 {A:1 B:1 C:5 E:1050 A=BS B=Tune C=BS D=BS 0R08=0.39 2R dump in. A C dump(s) in. Beam to tagger yoke} { } ALL

Comments:
List of off-loop magnets:

MBC2C19V.LOOP MBC2C20H.LOOP MBC2C21H.LOOP MBC2C21V.LOOP MBC2C22H.LOOP MBC2C23V.LOOP MBCAT01V.LOOP MBT1E01V.LOOP MBT1E02H.LOOP MBT2S02H.LOOP MBT2S05H.LOOP MBT2S06V.LOOP MBT2S10V.LOOP MCE3P03H.LOOP MQD0R03.LOOP MQE3M03.LOOP MQM1H02.LOOP MQO1H02.LOOP MQO1H03.LOOP MQO1H03A.LOOP

1st pass extraction check - looks OK

Log Number: 3819096Logbook(s): Entry Makers: eformanBody: 

Checking 1st pass extraction after unplanned down & work today. We're a little over separated at 17 mm but leaving it alone for now.

Extracted beam: fig. 2
Recirc beam: fig.5

FSD Masking Tool: Successful System Masking for: BeamLossMonitor

Log Number: 3819095Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: eforman
Areas: All

BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T09
ILM4C01
ILM4C02
ILM4C03
ILM1C0

Back in Business!

Log Number: 3819092Logbook(s): Entry Makers: aikenBody: 

Separators Restored. SL C100's restored without too much trouble. 1L07 took a lot of extra TLC, but all cavities are on now. Restoring beam to HB for steering and Harp Scans.

COMMENT ON Energizing and Commissioning BSY3C PS

Log Number: 3819087Logbook(s): Body: 

Summary from yesterday (7/31)

After the move into the building (ATLIS 20982 and 20980), relevant portions of Steps 1 through 8 were carried out by M.Todd and M.Kissinger. Interlock checks were carried out with help from O.Kumar.
Thanks to G.Croke for remotely configuring the controls software and ioc's on short noticed with interlock checks.

Pages